You are designing an experiment for a research project. A very delicate computer sensor requires that...

Question 1 Unsaved

You are designing an experiment for a research project. A very delicate computer sensor requires that it be operated at a voltage of 6 V and a power of 3.5 W. If the wires connecting the experiment to its power source have a total resistance (for all wires) of 0.7 Ω and the power supply has an interenal resistance of 0.9 Ω, what must you design the EMF of the power source to be?

Your answer should be in volts to 3 significant figures. Do not include units in your answer.
Answer #1

We know that

Power P =V*I

I =P/V =3.5/6 =0.583A

The load resistance is given by

RL =6V/(0.583A) =10.29ohms

what must you design the EMF of the power source to be V = (0.583)(0.9+0.7+10.29) =6.929V